# A Survey on Self-Evolution of Large Language Models

### v1

@@ 3.1.1 LLMs
Instruction Following: The capability to follow instructions is essential for effectively applying language models. It allows these models to address specific user needs across different tasks and domains, aligning their responses within the given context Xu et al. (2023a).
Reasoning: LLMs can self-evolve to recognize statistical patterns, making logical connections and deductions based on the information. They evolve to perform better reasoning involving methodically dissecting problems in a logical sequence Cui and Wang (2023).
Math: LLMs enhance the intricate ability to solve mathematical problems covering arithmetic, math word, geometry, and automated theorem proving Ahn et al. (2024) towards self-evolution.
Coding: Methods improve the LLM coding abilities to generate more precise and robust programs Singh et al. (2023); Zelikman et al. (2023). Furthermore, EvoCodeBench Li et al. (2024a) provides an evolving benchmark that updates periodically to prevent data leakage.
Role-Play: It involves an agent understanding and acting out a particular role within a given context. This is crucial in scenarios where the model must fit into a social structure or follow a set of behaviors associated with a specific identity or function Lu et al. (2024a).
Others: Apart from the above fundamental evolution objectives, self-evolution can also achieve and a wide range of NLP tasks Stammer et al. (2023); Koa et al. (2024); Gulcehre et al. (2023); Zhang et al. (2024b, c).
@@ 3.1.2 LLM-based Agents
Planning: It involves the ability to strategize and prepare for future actions or goals. An agent with this skill can analyze the current state, predict the outcomes of potential actions, and create a sequence of steps to achieve a specific objective Qiao et al. (2024).
Tool Use: This is the capacity to employ objects or instruments in the environment to perform tasks, manipulate surroundings, or solve problems Zhu et al. (2024).
Embodied Control: It refers to an agent’s ability to manage and coordinate its physical form within an environment. This encompasses locomotion, dexterity, and the manipulation of objects Bousmalis et al. (2023).
Communication: It is the skill to convey information and understand messages from other agents or humans. Agents with advanced communication abilities can participate in dialogue, collaborate with others, and adjust their behaviour based on the communication received Ulmer et al. (2024).
@@ 3.2 Evolution Directions
Improving Performance: The goal is to continuously enhance the model’s understanding and generation across various languages and abilities. For instance, a model initially trained for question answering and chitchat can autonomously extend its proficiency and develop abilities like diagnostic dialogue Tu et al. (2024), social skills Wang et al. (2024e), and role-playing Lu et al. (2024a).
@@
Adaptation to Feedback: This involves improving model responses based on feedback to better align with preferences or adapt to environments Yang et al. (2023a); Sun et al. (2024).
@@
Expansion of Knowledge Base: The aim is to continuously update the model’s knowledge base with the latest information and trends. For example, a model might automatically integrate new scientific research into its responses Wu et al. (2024).
@@
Safety, Ethic and Reducing Bias: The goal is to identify and mitigate biases in the model’s responses, ensuring fairness and safety. One effective strategy is to incorporate guidelines, such as constitutions or specific rules, to identify inappropriate or biased responses and corrected through model updates Bai et al. (2022); Lu et al. (2024b).
Task $f^{{\mathcal{T}}}$ Solution $f^{{\mathcal{Y}}}$ Feedback $f^{{\mathcal{F}}}$
Large Language Models
Self-Align Sun et al. (2024) Context-Based Pos-G - Filtering In-W IF
SciGLM Zhang et al. (2024b) Context-Based - - - In-W Other
EvIT Tao et al. (2024a) Context-Based - - - In-W Reasoning
MEEL Tao et al. (2024b) Context-Based - - - In-W Reasoning
UltraChat Ding et al. (2023) Context-Based - - - In-W Role-Play
SOLID Askari et al. (2024) Context-Based Pos-S - Filtering In-W Role-Play
Ditto Lu et al. (2024a) Context-Based Pos-S, Neg-P - - In-W Role-Play
MetaMath Yu et al. (2023b) Context-Free Pos-R - - In-W Math
Self-Rewarding Yuan et al. (2024) Context-Free - Model - In-W IF,Reasoning,Role-Play
Kun Zheng et al. (2024b) Context-Free - - Filtering In-W IF,Reasoning
PromptBreeder Fernando et al. (2023) Context-Free - - - In-C Math, Reasoning
Ada-Instruct Cui and Wang (2023) Context-Free - - - In-W Math, Reasoning, Code
Backtranslation Li et al. (2023b) Context-Free - - - In-W IF
DiverseEvol Wu et al. (2023) Selective Pos-I - - In-W Code
Grath Chen and Li (2024) Selective Neg-C Model - In-W Reasoning
RESTem Singh et al. (2023) Selective - Model Filtering In-W Math, Code
SOFT Wang et al. (2024c) Selective - - - In-W IF
LSX Stammer et al. (2023) - Pos-R Model Correcting In-W Other
LMSI Huang et al. (2022) - Pos-R - Filtering In-W Math
TRAN Yang et al. (2023b) - Pos-G - - In-C Reasoning
MOT Li and Qiu (2023) - Pos-R, Pos-G - Filtering In-C Math, Reasoning
STaR Zelikman et al. (2022) - Pos-R, Neg-C Model Correct In-W Reasoning
COTERRORSET Tong et al. (2024) - Pos-R, Neg-C - - In-W Math, Reasoning
Self-Debugging Chen et al. (2023c) - Pos-I Env - In-C Code
SelfEvolve Jiang et al. (2023) - Pos-I - - In-C Code
Reflexion Shinn et al. (2024) - Pos-I, Pos-G - - In-C Code, Reasoning
V-STaR Hosseini et al. (2024) - Neg-C Model Filter In-W Math, Code
Self-Contrast Zhang et al. (2024e) - Neg-C Model - In-W Reasoning
SALMON Sun et al. (2023) - Neg-C Model - In-W IF,Reasoning,Role-Play
SPIN Chen et al. (2024) - Neg-C - - In-W IF,Reasoning,Role-Play
RLCD Yang et al. (2023a) - Neg-P Model - In-W IF
DLMA Liu et al. (2024a) - Neg-P Model - In-W IF
SELF Lu et al. (2023) - - Model Correct In-W IF, Math
LLM Agents
AutoAct Qiao et al. (2024) Context-Based Pos-I Env Filtering In-W Planning, Tool
KnowAgent Zhu et al. (2024) Context-Based Pos-I, Pos-G Env Filtering In-W Embodied, Planning, Tool
RoboCat Bousmalis et al. (2023) Context-Free Pos-I Env - In-W Embodied
STE Wang et al. (2024b) Context-Free Pos-I, Neg-C Env Correct In-W Tool
IML Wang et al. (2024a) - Pos-R, Pos-G - - In-C Reasoning
SinViG Xu et al. (2024b) - Pos-I Env Filtering In-W Embodied
ETO Song et al. (2024) - Pos-I, Neg-C Env Correct In-W Tool
A3Yang et al. (2024c) - Pos-I, Neg-C Env Correct In-W Tool
Debates Taubenfeld et al. (2024) - Pos-S - - In-W Communication
SOTOPIA-$\pi$ Wang et al. (2024e) - Pos-S,Pos-G Env - In-W Communication
Self-Talk Ulmer et al. (2024) - Pos-S, Pos-G Model Filtering In-W Communication
MemGPT Packer et al. (2023) - Pos-G Env Filtering In-C Communication
MemoryBank Zhong et al. (2024b) - Pos-G Env Filtering In-C Communication
ProAgent Zhang et al. (2024a) - Pos-G Env - In-C Embodied
Agent-Pro Zhang et al. (2024d) - Pos-G Env - In-C Planning
AesopAgent Wang et al. (2024d) - Pos-G Env - In-C Planning
ICE Qian et al. (2024) - Pos-G Env - In-C Planning
TiM Liu et al. (2023a) - Pos-G - - In-C Communication
Werewolf Xu et al. (2023b) - Pos-G - - In-C Planning
Table 1: Overview of self-evolution methods, detailing approaches across evolutionary stages. Key: Pos (Positive), Neg (Negative), R (Rationale-based), I (Interactive), S (Self-play), G (Grounded), C (Contrastive), P (Perturbative), Env (Environment), In-W (In-Weight), In-C (In-Context), IF (Instruction-Following). For the evolution objectives, Adaptation to Feedback is in green, Expansion of Knowledge Base is in blue, and Safety, Ethic and Reducing Bias is in brown. Improving Performance is in the default color, black.
@@ 4 Experience Acquisition
Exploration and exploitation Gupta et al. (2006) are fundamental strategies for learning in humans and LLMs. Among that, exploration involves seeking new experiences to achieve objectives and is analogous to the initial phase of LLM self-evolution, known as experience acquisition. This process is crucial for self-evolution, enabling the model to autonomously tackle core challenges such as adapting to new tasks, overcoming knowledge limitations, and enhancing solution effectiveness. Furthermore, experience is a holistic construct, encompassing not only the tasks encountered Dewey (1938) but also the solutions developed to address these tasks Schön (2017), and the feedback Boud et al. (2013) received as a result of task performance.
@@ 4.1 Task Evolution
The first kind of knowledge is structured. Structured knowledge is dense in information and well-organized. Self-Align Sun et al. (2024) curates topic-guided tasks generated covering 20 scientific topics such as scientific and legal expertise. Apart from topic knowledge, DITTO Lu et al. (2024a) includes character knowledge from Wikidata and Wikipedia. The knowledge comprises attributes, profiles, and concise character details for role-play conversations. SOLID Askari et al. (2024) generates structured entity knowledge as conversation starters.
The second group consists of tasks evolving from an unstructured context. Unstructured context is easy to obtain but is sparse in knowledge. UltraChat Ding et al. (2023) gathers unstructured knowledge of 20 types of text materials based on 30 meta-concepts to construct conversation tasks. SciGLM Zhang et al. (2024b) derives questions from the text of diversified science subjects, which covers rich scientific knowledge. EvIT Tao et al. (2024a) derives event reasoning tasks based on large-scale unstructured events mined from the unsupervised corpus. Similarly, MEEL Tao et al. (2024b) evolves multi-modal events in both image and text to construct the tasks for MM event reasoning.
@@
First, the LLMs can prompt themselves to generate new tasks according to ${\mathcal{E}}^{t}$. Self-Instruct Wang et al. (2023b); Honovich et al. (2022); Roziere et al. (2023) is a typical methodology of Knowledge-Free task evolution. These methods self-generate a variety of new task instructions based on evolution objectives. Ada-Instruct Cui and Wang (2023) further proposes an adaptive task instruction generation strategy that fine-tunes open-source LLMs to generate lengthy and complex task instructions for code completion and mathematical reasoning.
Second, extending and boosting original tasks increases the quality of instructions. WizardLM Xu et al. (2023a) proposes Evol-Instruct that evolves instruction following tasks with in-depth and in-breadth evolving and further expands it in code generation Luo et al. (2024). MetaMath Yu et al. (2023b) rewrites the question in multiple ways, including rephrasing, self-verification, and FOBAR. It evolves a new MetaMathQA dataset for fine-tuning LLMs to improve mathematical task-solving. Promptbreeder Fernando et al. (2023) evolves seed tasks via mutation prompts. It further evolves mutation prompts via the hyper mutation prompts to increase the task diversity.
Third, deriving tasks from plain text is another way. Backtranslation Li et al. (2023b) extracts self-contained segments in unlabelled data and regards it as the answers to tasks. Similarly, Kun Zheng et al. (2024b) presents a task self-evolving algorithm utilizing instruction harnessed from unlabelled data towards back-translation.
@@
Instead of task generation, we may start with large-scale existing tasks. At each iteration, LLMs can select tasks that exhibit the highest relevance to the current evolving objective $\mathcal{E}^{t}$ without additional generation. This approach obviates the intricate curation of new tasks, streamlining the evolution process Zhou et al. (2024); Li et al. (2023a); Chen et al. (2023a).
A simple task selecting method is to randomly sample tasks from the task pool like REST Gulcehre et al. (2023), RESTem Singh et al. (2023), and GRATH Chen and Li (2024) do. Rather than random selection, DIVERSE-EVOL Wu et al. (2023) introduces a data sampling technique where the model selects new data points based on their distinctiveness in the embedding space, ensuring diversity enhancement in the chosen subset. SOFT Wang et al. (2024c) then splits the initial training set. Each iteration selects one chunk of the split set as the evolving task.
Li et al. (2024b) propose Selective Reflection-Tuning and select a subset of tasks via a novel metric calculating to what extent the answer is related to the question. V-STaR Hosseini et al. (2024) selects the correct solutions in the previous iteration and adds their task instructions to the task set of the next iteration.
@@ 4.2 Solution Evolution
After obtaining evolved tasks, LLMs solve the tasks to acquire the corresponding solution. The most common strategy is to generate the solution directly according to the task formulation Zelikman et al. (2022); Gulcehre et al. (2023); Singh et al. (2023); Zheng et al. (2024b); Yuan et al. (2024). However, this straightforward approach might reach solutions irrelevant to the evolution objective, leading to suboptimal evolution Hare (2019). Therefore, solution evolution uses different strategies to solve tasks and enhance LLM capabilities by ensuring that solutions are not just generated but are also relevant and informative. In this section, we comprehensively survey these strategies and illustrate them in Figure 5. We first formulate the solution-evolution as follows:
@@ 4.2.1 Positive
The model incorporates rationale explanations towards approaching the evolving objective when solving the tasks and can self-evolve by utilizing such rationales. These methods enable models to explicitly acknowledge the evolution objective and complete this task in that direction Wei et al. (2022); Yao et al. (2024); Besta et al. (2024); Yao et al. (2022).
Huang et al. (2022) proposes a method where an LLM self-evolves using "high-confidence" rationale-augmented answers generated for unlabeled questions. Similarly, STaR Zelikman et al. (2022) generates rationale when solving the task. If the answer is wrong, it further corrects the rationale and the answer. Then, it uses the answer and rationale as experiences to fine-tune the model. Similarly, LSX Stammer et al. (2023) proposes the novel paradigm to generate an explanation of the answer, incorporating an iterative loop between a learner module performing a base task and a critic module that assesses the quality of explanations given by the learner. Song et al. (2024); Yang et al. (2024c) obtain rationales in the ReAct Yao et al. (2022) style when solving the tasks. The rationales are further engaged in training the agents in the following step.
@@
SelfEvolve and LDB Jiang et al. (2023); Zhong et al. (2024a) improve code generation ability via self-evolution. They allow the model to generate code and acquire feedback via running the code on the interpreter. As another environment, Song et al. (2024); Yang et al. (2024c) interact in embodied scenarios and acquire feedback. They learn to take proper actions based on their current state. For agent abilities, AutoAct Qiao et al. (2024) introduces self-planning from scratch, focusing on an intrinsic self-learning process. In this process, agents enhance their abilities through recursive planning iterations with environment feedback. Following AutoAct, Zhu et al. (2024) further enhances agent training by integrating self-evolution and an external action knowledge base. This approach guides action generation and boosts planning ability through environment-driven corrective feedback loops.
@@
It’s the situation where a model learns to evolve by playing against copies of itself. Self-play is a powerful evolving method because it enables systems to communicate with themselves to get feedback in a closed loop. It’s especially effective in environments where the model can simulate various sides of the roles, like multi-player games Silver et al. (2016, 2017). Compared with interactive methods, self-play is an effective strategy to obtain feedback without an environment.
Taubenfeld et al. (2024) investigate the systematic biases in simulations of debates by LLMs. On the contrary to debating, Ulmer et al. (2024) engage LLMs in conversations following generated principles. Another kind of conversation via role-playing. Lu et al. (2024a) proposes self-simulated role-play conversation. The process involves instructing the LLM with character profiles and aligning its responses to maintain consistency with the character’s knowledge and style. Similarly, Askari et al. (2024) propose SOLID to generate large-scale intent-aware role-play dialogues. This self-playing aspect harnesses the expansive knowledge of LLMs to construct information-rich exchanges that streamline the dialog generation process. Wang et al. (2024e) introduces a novel approach whereby each LLM follows a role and communicates with each other to achieve their goals.
@@
To reach the evolving objective and reduce exploration space, models can be grounded on existing rules Sun et al. (2024) and previous experiences for further explicit guidance when solving the tasks.
LLMs can generate desirable solutions more effectively by being grounded on pre-defined rules and principles. For instance, Self-Align Sun et al. (2024) generated self-evolved questions with principle-driven constraints to guide the task-solving process. SALMON Sun et al. (2023) design a set of combined principles that requires the model to follow when solving the task. Self-Talk Ulmer et al. (2024) ensures the LLMs generate a workflow-aligned conversation based on preset agent characters. They generate the workflow in advance based on GPT-4.
Besides pre-defined rules, grounding on previous experiences can improve the solutions. MemoryBank Zhong et al. (2024b) and TiM Liu et al. (2023a) answer current questions by incorporating previous question-answer records. Rather than previous solution histories, MoT Li and Qiu (2023), IML Wang et al. (2024a), and TRAN Yang et al. (2023b) incorporate induced rules from the histories to answer new questions. MemGPT Packer et al. (2023) combines these merits and retrieves previous questions, solutions, induced events, and user portrait knowledge.

### v2

@@ 3.1.1 LLMs
Instruction Following: The capability to follow instructions is essential for effectively applying language models. It allows these models to address specific user needs across different tasks and domains, aligning their responses within the given context [134].
Reasoning: LLMs can self-evolve to recognize statistical patterns, making logical connections and deductions based on the information. They evolve to perform better reasoning involving methodically dissecting problems in a logical sequence [27].
Math: LLMs enhance the intricate ability to solve mathematical problems covering arithmetic, math word, geometry, and automated theorem proving [2] towards self-evolution.
Coding: Methods improve the LLM coding abilities to generate more precise and robust programs [103, 150]. Furthermore, EvoCodeBench [62] provides an evolving benchmark that updates periodically to prevent data leakage.
Role-Play: It involves an agent understanding and acting out a particular role within a given context. This is crucial in scenarios where the model must fit into a social structure or follow a set of behaviors associated with a specific identity or function [77].
Others: Apart from the above fundamental evolution objectives, self-evolution can also achieve and a wide range of NLP tasks [106, 59, 42, 152, 153].
@@ 3.1.2 LLM-based Agents
Planning: It involves the ability to strategize and prepare for future actions or goals. An agent with this skill can analyze the current state, predict the outcomes of potential actions, and create a sequence of steps to achieve a specific objective [90].
Tool Use: This is the capacity to employ objects or instruments in the environment to perform tasks, manipulate surroundings, or solve problems [165].
Embodied Control: It refers to an agent’s ability to manage and coordinate its physical form within an environment. This encompasses locomotion, dexterity, and the manipulation of objects [13].
Communication: It is the skill to convey information and understand messages from other agents or humans. Agents with advanced communication abilities can participate in dialogue, collaborate with others, and adjust their behaviour based on the communication received [118].
@@ 3.2 Evolution Directions
Improving Performance: The goal is to continuously enhance the model’s understanding and generation across various languages and abilities. For instance, a model initially trained for question answering and chitchat can autonomously extend its proficiency and develop abilities like diagnostic dialogue [117], social skills [125], and role-playing [77].
@@
Adaptation to Feedback: This involves improving model responses based on feedback to better align with preferences or adapt to environments [138, 108].
@@
Expansion of Knowledge Base: The aim is to continuously update the model’s knowledge base with the latest information and trends. For example, a model might automatically integrate new scientific research into its responses [132].
@@
Safety, Ethic and Reducing Bias: The goal is to identify and mitigate biases in the model’s responses, ensuring fairness and safety. One effective strategy is to incorporate guidelines, such as constitutions or specific rules, to identify inappropriate or biased responses and correct them through model updates [9, 78].
@@ 4 Experience Acquisition
Exploration and exploitation [44] are fundamental strategies for learning in humans and LLMs. Among that, exploration involves seeking new experiences to achieve objectives and is analogous to the initial phase of LLM self-evolution, known as experience acquisition. This process is crucial for self-evolution, enabling the model to autonomously tackle core challenges such as adapting to new tasks, overcoming knowledge limitations, and enhancing solution effectiveness. Furthermore, experience is a holistic construct, encompassing not only the tasks encountered [31] but also the solutions developed to address these tasks [96], and the feedback [12] received as a result of task performance.
@@ 4.1 Task Evolution
The first kind of knowledge is structured. Structured knowledge is dense in information and well-organized. Self-Align [108] guides task generation by covering 20 topics, such as scientific and legal expertise, to ensure diversity. Apart from topic knowledge, DITTO [77] includes character knowledge from Wikidata and Wikipedia. The knowledge comprises attributes, profiles, and concise character details for role-play conversations. SOLID [6] generates structured entity knowledge as conversation starters.
The second group consists of tasks evolving from an unstructured context. Unstructured context is easy to obtain but is sparse in knowledge. UltraChat [32] gathers unstructured knowledge of 20 types of text materials to construct questions or instructions. SciGLM [152] derives questions from the text of diversified science subjects, which covers rich scientific knowledge. EvIT [110] derives event reasoning tasks based on large-scale unstructured events mined from the unsupervised corpus. Similarly, MEEL [111] evolves multi-modal events in both image and text to construct the tasks for MM event reasoning.
@@
First, the LLMs can prompt themselves to generate new tasks according to ${\mathcal{E}}^{t}$. Self-Instruct [127, 48, 93] is a typical methodology of Knowledge-Free task evolution. These methods self-generate a variety of new task instructions based on evolution objectives. Ada-Instruct [27] further proposes an adaptive task instruction generation strategy that fine-tunes open-source LLMs to generate lengthy and complex task instructions for code completion and mathematical reasoning.
Second, extending and boosting original tasks increases the quality of instructions. WizardLM [134] proposes Evol-Instruct that evolves instruction following tasks with in-depth and in-breadth evolving and further expands it in code generation [79]. MetaMath [146] rewrites the question in multiple ways, including rephrasing, self-verification, and FOBAR. It evolves a new MetaMathQA dataset for fine-tuning LLMs to improve mathematical task-solving. Promptbreeder [36] evolves seed tasks via mutation prompts. It further evolves mutation prompts via the hyper mutation prompts to increase the task diversity.
Third, deriving tasks from plain text is another way. Backtranslation [67] extracts self-contained segments in unlabelled data and regards it as the answers to tasks. Similarly, Kun [158] presents a task self-evolving algorithm utilizing instruction harnessed from unlabelled data towards back-translation.
@@
Instead of task generation, we may start with large-scale existing tasks. At each iteration, LLMs can select tasks that exhibit the highest relevance to the current evolving objective $\mathcal{E}^{t}$ without additional generation. This approach obviates the intricate curation of new tasks, streamlining the evolution process [162, 63, 19].
A simple task selecting method is to randomly sample tasks from the task pool like REST [42], RESTem [103], and GRATH [21] do. Rather than random selection, DIVERSE-EVOL [131] introduces a data sampling technique where the model selects new data points based on their distinctiveness in the embedding space, ensuring diversity enhancement in the chosen subset. SOFT [122] then splits the initial training set. Each iteration selects one chunk of the split set as the evolving task.
Li et al [64] propose Selective Reflection-Tuning and select a subset of tasks via a novel metric calculating to what extent the answer is related to the question. V-STaR [49] selects the correct solutions in the previous iteration and adds their task instructions to the task set of the next iteration.
@@ 4.2 Solution Evolution
After obtaining evolved tasks, LLMs solve the tasks to acquire the corresponding solution. The most common strategy is to generate the solution directly according to the task formulation [149, 42, 103, 158, 148]. However, this straightforward approach might reach solutions irrelevant to the evolution objective, leading to suboptimal evolution [45]. Therefore, solution evolution uses different strategies to solve tasks and enhance LLM capabilities by ensuring that solutions are not just generated but are also relevant and informative. In this section, we comprehensively survey these strategies and illustrate them in Figure 5. We first formulate the solution-evolution as follows:
@@ 4.2.1 Positive
The model incorporates rationale explanations towards approaching the evolving objective when solving the tasks and can self-evolve by utilizing such rationales. These methods enable models to explicitly acknowledge the evolution objective and complete this task in that direction [128, 145, 10, 144].
Huang et al [52] proposes a method where an LLM self-evolves using ”high-confidence” rationale-augmented answers generated for unlabeled questions. Similarly, STaR [149] generates rationale when solving the task. If the answer is wrong, it further corrects the rationale and the answer. Then, it uses the answer and rationale as experiences to fine-tune the model. Similarly, LSX [106] proposes the novel paradigm to generate an explanation of the answer, incorporating an iterative loop between a learner module performing a base task and a critic module that assesses the quality of explanations given by the learner. Song et al [105], Yang et al [141] obtain rationales in the ReAct [144] style when solving the tasks. The rationales are further engaged in training the agents in the following step.
@@
SelfEvolve and LDB [56, 159] improve code generation ability via self-evolution. They allow the model to generate code and acquire feedback via running the code on the interpreter. As another environment, Song et al [105], Yang et al [141] interact in embodied scenarios and acquire feedback. They learn to take proper actions based on their current state. For agent abilities, AutoAct [90] introduces self-planning from scratch, focusing on an intrinsic self-learning process. In this process, agents enhance their abilities through recursive planning iterations with environment feedback. Following AutoAct, [165] further enhances agent training by integrating self-evolution and an external action knowledge base. This approach guides action generation and boosts planning ability through environment-driven corrective feedback loops.
@@
It’s the situation where a model learns to evolve by playing against copies of itself. Self-play is a powerful evolving method because it enables systems to communicate with themselves to get feedback in a closed loop. It’s especially effective in environments where the model can simulate various sides of the roles, like multi-player games [101, 102]. Compared with interactive methods, self-play is an effective strategy to obtain feedback without an environment.
Taubenfeld et al [112] investigate the systematic biases in simulations of debates by LLMs. On the contrary to debating, Ulmer et al [118] engage LLMs in conversations following generated principles. Another kind of conversation via role-playing. Lu et al [77] proposes self-simulated role-play conversation. The process involves instructing the LLM with character profiles and aligning its responses to maintain consistency with the character’s knowledge and style. Similarly, Askari et al [6] propose SOLID to generate large-scale intent-aware role-play dialogues. This self-playing aspect harnesses the expansive knowledge of LLMs to construct information-rich exchanges that streamline the dialog generation process. Wang et al [125] introduces a novel approach whereby each LLM follows a role and communicates with each other to achieve their goals.
@@
To reach the evolving objective and reduce exploration space, models can be grounded on existing rules [108] and previous experiences for further explicit guidance when solving the tasks.
LLMs can generate desirable solutions more effectively by being grounded on pre-defined rules and principles. For instance, Self-Align [108] generated self-evolved questions with principle-driven constraints to guide the task-solving process. SALMON [107] design a set of combined principles that requires the model to follow when solving the task. Self-Talk [118] ensures the LLMs generate a workflow-aligned conversation based on preset agent characters. They generate the workflow in advance based on GPT-4.
Besides pre-defined rules, grounding on previous experiences can improve the solutions. MemoryBank [161] and TiM [72] answer current questions by incorporating previous question-answer records. Rather than previous solution histories, MoT [66], IML [121], and TRAN [140] incorporate induced rules from the histories to answer new questions. MemGPT [84] combines these merits and retrieves previous questions, solutions, induced events, and user portrait knowledge.